Solve the given initial-value problem. x=(-1-2)x+(). xco) - -() X(t) =
Answer #1

soln :- x= ( 113 -2 J**(5) XO) to -6 7 1 3 3 4 Charceteastic en is -1 -2 1: 4-7 (x+1) (A-4) +6 EO 123*+2-0 (x-2)-1) - Eigenet debut Now to determine Xp X et 3 et X4 gelt geat - 082 t re et -t -ze ( -let eat eat xt (5) ) = 6 -t -ze pet det eat 16)-2540 1 25 +15 Xp = -35 40 Hence X() Xet Xp t -1 2+/-2 X(t) = de 3 But given XO) .6 7 -6 - G 22 35 teo e° f 7 362 (uo -G-202-